find the Expectation value of. [xp]-[px] indetail


1
Expert's answer
2021-02-22T10:25:51-0500

Answer

Suppose we have any wavefunction

"\psi" .

Now commutator

"[x, p_x]\psi=xp_x\psi-p_xx\psi"

"=x(-i\hbar \frac{d\psi}{dx}) -(-i\hbar\frac{d(\psi x) }{dx})"

"=-i\hbar x\frac{d\psi}{dx}+i\hbar x\frac{d\psi }{dx}+i\hbar\psi"

"=i\hbar"